# KIAS-Q25002

### v1

@@ 1 Introduction
in terms of the VEV $v=2\mu/\sqrt{\lambda}$, where $m=ev$ and $m_{h}=\lambda^{1/2}v/\sqrt{2}$ are the masses of the gauge and Higgs bosons. This Lagrangian is invariant under the gauge transformation $A^{\mu}\rightarrow A^{\mu}+\partial^{\mu}\Lambda$ and $\pi\rightarrow m+\Lambda$. If we regard the Goldstone field $\pi$ as the fifth gauge-field component $A^{4}$, we can obtain
@@ 3 Connecting the 5V $R_{\xi}$ gauge with other gauges
for the interactions, $AAh$, $AAhh$, $hhh$, and $hhhh$, of the gauge and Higgs bosons, $A$ and $h$, where the lower-index fifth components are defined as $P_{1,4}=g_{4M}P^{M}_{1}$ and $P_{2,4}=g_{4M}P^{M}_{2}$, from the five-momenta, $P_{1}$ and $P_{2}$, respectively. These interaction vertices are divergence-free:
@@ 4 Identification of the high-energy divergence in tree-level amplitudes
$+\bigg{(}e^{2}\epsilon\,h+\frac{e^{2}}{2}h^{2}\bigg{)}\bigg{(}A^{\mu}-\frac{1}{e\epsilon}\partial^{\mu}A^{4}\bigg{)}^{2}-4\eta^{2}\epsilon\,h^{3}-\eta^{2}h^{4},$ (49)
@@
in the Lagrangian (2), with the redefined Higgs field $\boldsymbol{H}$ and fifth gauge-field component $\boldsymbol{A}^{4}$. The Higgs and fifth-component fields in the polar basis $\Phi=\frac{1}{\sqrt{2}}(v+h)e^{iA^{4}/v}$ can be expressed by the combinations of $\boldsymbol{H}$ and $\boldsymbol{A}^{4}$ as
@@
$\mathcal{L}(x)=$ $-\frac{1}{4}(\boldsymbol{A}^{MN})^{2}+\frac{1}{2}(\partial^{\mu}\boldsymbol{h})^{2}-\frac{1}{2}m_{h}^{2}\boldsymbol{h}^{2}$
@@
where the 4V components $\boldsymbol{A}^{\mu}$ of the gauge field is equivalent to the polar-basis ones $A^{\mu}$. The BRST invariance of the Lagrangian ($\mbox{\bf s}\mathcal{L}(x)=0$) can be constructed by defining the BRST transformation of each field as
@@
$=\Gamma^{[\boldsymbol{A}\boldsymbol{A}\boldsymbol{h}]}_{\mu_{2}\sigma}(P_{2},Q)\,\bigg{(}\frac{(Q^{\sigma}+P_{2}^{\sigma}/2)(Q^{\rho*}-P_{1}^{\rho*})}{m^{2}}\bigg{)}\,\Gamma^{[\boldsymbol{A}\boldsymbol{A}\boldsymbol{h}]}_{\rho\mu_{1}}(-Q^{*},P_{1}),$ (70)
@@ Acknowledgments
We thank Prof. Andrea Wulzer and Prof. Alex Pomarol for their valuable comments on this work. This work was supported by a KIAS Individual Grant (QP090001) through the Quantum Universe Center at the Korea Institute for Advanced Study. We also acknowledge the Institut de Física d’Altes Energies (IFAE) for their hospitality during our visit.

### v2

@@ 1 Introduction
in terms of the VEV $v=2\mu/\sqrt{\lambda}$, where $m=ev$ and $m_{h}=\lambda^{1/2}v/\sqrt{2}$ are the masses of the gauge and Higgs bosons. This Lagrangian is invariant under the gauge transformation $A^{\mu}\rightarrow A^{\mu}+\partial^{\mu}\Lambda$ and $\pi\rightarrow\pi+m\Lambda$. If we regard the Goldstone field $\pi$ as the fifth gauge-field component $A^{4}$, we can obtain
@@ 3 Connecting the 5V $R_{\xi}$ gauge with other gauges
for the interactions, $AAh$, $AAhh$, $hhh$, and $hhhh$, of the gauge and Higgs bosons, $A$ and $h$, where the lower-index fifth components are defined as $P_{1,4}=g_{4M}P^{M}_{1}$ and $P_{2,4}=g_{4M}P^{M}_{2}$, from the five-momenta, $P_{1}$ and $P_{2}$, respectively. Each of the simplified matrix expressions in Eqs. (36) and (37) represents a 4 by 4 matrix in the top left, four-component column and row vectors in the top right and bottom left, and a scalar element in the bottom right. The interaction vertices are divergence-free:
@@ 4 Identification of the high-energy divergence in tree-level amplitudes
$+\bigg{(}e^{2}\epsilon\,h+\frac{e^{2}}{2}h^{2}\bigg{)}\bigg{(}A^{\mu}-\frac{1}{e\epsilon}\partial^{\mu}A^{4}\bigg{)}^{2}-\frac{\lambda}{4}\epsilon\,h^{3}-\frac{\lambda}{16}h^{4},$ (49)
@@
in the Lagrangian (2), with the redefined Higgs field $\boldsymbol{H}$ and fifth gauge-field component $\boldsymbol{A}^{4}$. The Higgs and fifth-component fields in the polar basis $\Phi=\frac{1}{\sqrt{2}}He^{iA^{4}/v}$ with $H=v+h$ can be expressed by the combinations of $\boldsymbol{H}$ and $\boldsymbol{A}^{4}$ as
@@
$\mathcal{L}(x)=$ $-\frac{1}{4}(\boldsymbol{F}^{MN})^{2}+\frac{1}{2}(\partial^{\mu}\boldsymbol{h})^{2}-\frac{1}{2}m_{h}^{2}\boldsymbol{h}^{2}$
@@
with $\boldsymbol{F}^{MN}=\partial^{M}\boldsymbol{A}^{N}-\partial^{N}\boldsymbol{A}^{M}$ where the 4V components $\boldsymbol{A}^{\mu}$ of the gauge field is equivalent to the polar-basis ones $A^{\mu}$. The BRST invariance of the Lagrangian ($\mbox{\bf s}\mathcal{L}(x)=0$) can be constructed by defining the BRST transformation of each field as
@@
$=\Gamma^{[\boldsymbol{A}\boldsymbol{A}\boldsymbol{h}]}_{\mu_{2}\sigma}(P_{2},Q)\,\bigg{(}\frac{(Q^{\sigma}+P_{2}^{\sigma}/2)(Q^{\rho*}-P_{1}^{\rho*}/2)}{m^{2}}\bigg{)}\,\Gamma^{[\boldsymbol{A}\boldsymbol{A}\boldsymbol{h}]}_{\rho\mu_{1}}(-Q^{*},P_{1}),$ (70)
@@ Acknowledgments
We thank Prof. Andrea Wulzer, Prof. Alex Pomarol, and Dr. Kiyoharu Kawana for their valuable comments on this work. This work was supported by a KIAS Individual Grant (QP090001) through the Quantum Universe Center at the Korea Institute for Advanced Study. We also acknowledge the Institut de Física d’Altes Energies (IFAE) for their hospitality during our visit.